OpenStudy (anonymous):

What is the y coordinate of the y-intercept of the line that passes through the points (-4,-1) and (4,5)?

OpenStudy (anonymous):

wot's the firest step ?,! dunow it ?

OpenStudy (anonymous):

This is all the question gives me to answer. Sorry

OpenStudy (anonymous):

standard equation is y = m x + c m = slope = (y2-y1)/(x2-x1)= (5-(-1))/(4-(-4))=6/8=3/4 so equation obtained is y = (3/4)x+c Now to obtain c lets put x= 4 and y =5 (given in question) in above equation y = (3/4)x+c=>5=(3/4)4+c => c=5-3=2 so equation of line is y = (3/4)x+2 now line will intercept at y. the corresponding value of x for which the intercept will occur is x = 0. so lets put x = 0 in the equation we just obtained y = (3/4)x+2=> y = (3/4)*0+2=> y = 2 line will intersect y axis at y = 2 units

OpenStudy (anonymous):

line will intersect y axis at y=2 units brahhh

OpenStudy (anonymous):

Thank you!